OpenStudy (anonymous):

rewrite with a single exponent (9^7)^3

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <--- use the 3rd rule listed there

geerky42 (geerky42):

\((9^7)^3=(9^7)(9^7)(9^7)=9^{7+7+7}=9^{3\cdot7}=?\)